OpenStudy (anonymous):

Evaluate the function f(x) = 2x + 2, when x = –1. A. 0 B. 3 C. –4 D. 4

OpenStudy (texaschic101):

basically, all you have to do is sub in -1 for x f(x) = 2x + 2 f(-1) = 2(-1) + 2 f(-1) = -2 + 2 f(-1) = 0

OpenStudy (anonymous):

thts wat i thought just wanted to make sure

OpenStudy (texaschic101):

its always good to check if your not sure...confirmation is good :)

OpenStudy (anonymous):

thaxs

OpenStudy (texaschic101):

sure thing :)